# Discrete Distribution Networks

### v2

@@ Abstract
We introduce a novel generative model, the Discrete Distribution Networks (DDN), that approximates data distribution using hierarchical discrete distributions. We posit that since the features within a network inherently capture distributional information, enabling the network to generate multiple samples simultaneously, rather than a single output, may offer an effective way to represent distributions. Therefore, DDN fits the target distribution, including continuous ones, by generating multiple discrete sample points. To capture finer details of the target data, DDN selects the output that is closest to the Ground Truth (GT) from the coarse results generated in the first layer. This selected output is then fed back into the network as a condition for the second layer, thereby generating new outputs more similar to the GT. As the number of DDN layers increases, the representational space of the outputs expands exponentially, and the generated samples become increasingly similar to the GT. This hierarchical output pattern of discrete distributions endows DDN with unique property: more general zero-shot conditional generation. We demonstrate the efficacy of DDN and its intriguing properties through experiments on CIFAR-10 and FFHQ. The code is available at https://discrete-distribution-networks.github.io/
@@ 1 Introduction
To accurately fit the target distribution of large datasets, a substantial representational space is required. In the most extreme scenario, this space must be larger than the number of training data samples. However, current neural networks lack the feasibility to generate such a vast number of samples simultaneously. Therefore, we adopt a strategy from autoregressive models Van Den Oord et al. (2016) and partition this large space into a hierarchical conditional probability model. Each layer of this model needs only a small number of outputs. We then select one of these outputs as the output for that layer and use it as conditional input to the next layer. As a result, the output of the next layer will be more closely related to the selected conditional sample. If the number of layers is $L$ and the number of outputs per layer is $K$, then the output space of the network is $K^{L}$. Due to its exponential nature, this output space will be much larger than the number of samples in the dataset. fig. 1 shows how DDN generates images.
We posit that the contributions of this paper are as follows: (1) We introduce a novel generative model, termed Discrete Distribution Networks (DDN), which exhibit a more straightforward and streamlined principle and form. (2) For training the DDN, we propose the “Split-and-Prune” optimization algorithm, and a range of practical techniques. (3) We conduct preliminary experiments and analysis on the DDN, showcasing its intriguing properties and capabilities, such as zero-shot conditional generation and highly compact representations.
@@ 2 Related Work
Deep Generative Model. Generative Adversarial Networks (GANs) Radford et al. (2016); Brock et al. (2019) and Variational Autoencoders (VAEs) Kingma & Welling (2014); van den Oord et al. (2018) are two early successful generative models. GANs reduce the divergence between the generated sample distribution and the target distribution through a game between the generator and discriminator. However, regular GANs cannot map samples back to the latent space, thus they cannot reconstruct samples. VAEs encode data into a simple distribution’s latent space through an Encoder, and the Decoder is trained to reconstruct the original data from this simple distribution’s latent space. Notably, VQ-VAE van den Oord et al. (2018), an important improvement of VAEs, achieves a more compact discrete representation by replacing the VAE’s latent space with discrete codebooks. However, this discrete representation still has a 2D structure, leading to a significant amount of redundant information. Moreover, VQ-VAE relies on a prior network to perform generative tasks. Autoregressive models van den Oord et al. (2016) , with their simple principles and methods, model the target distribution by decomposing the target data into conditional probability distributions of each component. They can also compute the exact likelihood of target samples. However, the efficiency of these models is reduced when dealing with image data, which is not suitable to be decomposed into a sequence of components. Normalizing flow Kingma & Dhariwal (2018) is another class of generative models that can compute the likelihood. They use invertible networks to construct a mapping from samples to a noise space, and during the generation stage, they map back from noise to samples. Energy Based Models (EBM) Goyal et al. (2021); Song et al. (2021) with their high-quality and rich generative results, have led to the rise of diffusion models. However, their multi-step iterative generation process requires substantial computational resources. The recent Idempotent Generative Network Shocher et al. (2024) introduces a novel approach by training a neural network to be idempotent, mapping any input to the target distribution effectively.
@@ 3 Discrete Distribution Networks
For computational efficiency, we adopted a decoder structure similar to the generator in GANs for coarse-to-fine image generation, as shown in fig. 3(b). We refer to this as the Single Shot Generator. As each layer of DDN can naturally input and output RGB domain data, DDN seamlessly support the Recurrence Iteration Paradigm fig. 3(c).
@@
How to generation. When the network performs the generation task, replacing the Guided Sampler with a random choice enables image generation. Given the exponential representational space of $K^{L}$ sample space and the limited number of samples in the training set, the probability of sampling an image with the same latent space as those in the training set is also exponentially low. For image reconstruction tasks, the process is almost identical to the training process, only substituting the training image with the target image to be reconstructed and omitting the L2 Loss part from the training process. The Final Generated Image in fig. 3(a) represents the final reconstruction result. The indices of the selected samples along the way form the latent representation of the target image $\mathbf{k}^{*}_{1:L}$, same as eq. 3. Therefore, the latent $\mathbf{k}^{*}_{1:L}$ is a sequence of integers with length $L$, which we regard as the hierarchical discrete representation of the target sample. The latent space exhibits a tree structure with $L$ layers and $K$ degrees per node, where each leaf node represents a sample space, and its latent representation indicates the indices of all nodes along the path to this leaf node, as shown in fig. 1(b). In the latent sequence, values placed earlier correspond to higher-level nodes in the tree, controlling the low-frequency information of the output sample, while later values tend to affect high-frequency information.
@@ 3.1 Optimization with Split-and-Prune
We have observed two primary issues resulting from each layer calculating loss only with the closest output samples to the ground truth (GT). Firstly, a problem similar to the“dead codebooks” in VQ-VAE arises, wherein output nodes that are not selected for a long time receive no gradient updates. During generation, these “dead nodes” are selected with equal probability, leading to poor output. The second issue is the probability density shift. For instance, in a one-dimensional asymmetric bimodal distribution, the target distribution is a mixture of two thin and tall Gaussian distribution functions, with one larger and one smaller peak. The means of these two peaks are -1 and 1, respectively. Therefore, half of the output samples with initial values less than 0 will be matched with samples from the larger peak and optimized towards the larger Gaussian distribution. Meanwhile, the other half of the output samples with values greater than 0 will be optimized towards the smaller Gaussian. A problem arises as the large and small peaks carry different probability masses but occupy equal portions of the output samples, resulting in the same sampling probability during generation.
Figure 5: Toy examples for two-dimensional data generation. The numerical values at the bottom of each figure represent the Kullback-Leibler (KL) divergence. Due to phenomena such as “dead nodes” and “density shift”, the application of Gradient Descent alone fails to properly fit the Ground Truth (GT) density. However, by employing the Split-and-Prune strategy, the KL divergence is reduced to even lower than that of the Real Samples. More details for this experiment are in the Appendix.
@@
Inspired by the theory of evolution, we propose a new optimization algorithm—Split-and-Prune—to address the above issues, as outlined in algorithm 1. The Split operation targets nodes frequently matched by training samples, while the Prune operation addresses the issue of “dead nodes”. These nodes are akin to species in evolution, subject to diversification and extinction. During training, we track the frequency with which each node is matched by training samples. For nodes with excessive frequency, we execute the Split operation, cloning the node into two new nodes, each inheriting half of the old node’s frequency. Although these two new nodes have identical parameters and outputs, the next matched training sample will only be associated with one node. Therefore, the loss and gradient only affect one node’s parameters. Consequently, their parameters and outputs exhibit slight differences, dividing the old node’s match space into two. In subsequent training, the outputs of the two new nodes will move towards the centers of their respective spaces under the influence of the loss, diverging to produce more diverse outputs. For nodes with low matching frequency (dead nodes), we implement the Prune operation, removing them outright. fig. 4 illustrates the process of the Split-and-Prune operation and how it reduces the distance between the discrete distribution represented by the Output Nodes and the target distribution. The efficacy of the Split-and-Prune optimization algorithm is validated through examples of fitting 2d density maps in fig. 5.
@@ 3.2 Applications
Zero-Shot Conditional Generation (ZSCG). Each layer of the DDN produces multiple target samples, with a selected sample being forwarded as a condition to the next layer. This enables the generation of new samples in the desired direction, ultimately producing a sample that meets the given condition. Indeed, the reconstruction process shown in fig. 1(a) is a ZSCG process guided by a target image, which never directly enters the network.
To implement ZSCG, we replace the Guided Sampler in fig. 3(a) with a Conditional Guided Sampler. For instance, when generating an image of class $y_{i}$ guided by a classifier $g_{cls}$, we replace the “L2 Distance to Training Image” in the Guided Sampler of fig. 3(a) with the probability of each output image belonging to class $y_{i}$ according to the classifier. We then replace “argmin” with “argmax” to construct the Guided Sampler for this classifier. Similar to eq. 1, the sampling method is as follow:
@@
After performing $L$ steps guided sampling and $L\times K$ steps of classification, the ZSCG results can be obtained without any gradient.
For super-resolution and colorization tasks, we construct a transform that converts the generated images to the target domain (low-resolution or grayscale). This approach significantly reduces the impact of the missing signal from the condition on the generated images, allowing DDN to successfully perform super-resolution tasks even when the source image has a resolution as low as $4\times 4$.
The use of “argmax” in the Guided Sampler causes each layer to select a fixed sample, resulting in a single image output for each condition, similar to the greedy sampling in GPT. To increase diversity, we employ more flexible sampling methods. For most ZSCG tasks, we use Top-k sampling with $k=2$, balancing diversity and condition appropriateness in the large generation space ($2^{L}$).
The versatility of ZSCG can be further enhanced by combining different Guided Samplers. For example, an image can guide the primary structure while text guides the attributes. The influence of each condition can be adjusted by setting their respective weights. Experiments involving the combination of different Guided Samplers will be presented in the Appendix.
Table 1: Quantitative comparison on CIFAR-10
Method Type FID↓
DCGAN GAN 37.1
IGEBM EBM 38.0
VAE VAE 106.7
Gated PixelCNN AR 65.9
GLOW Flow 46.0
DDN(ours) DDN 52.0
@@
In our experiments, we set $K$=512 as the default, considering the balance between generation performance and training efficiency. However, from the perspective of data compression, setting $K$ to 2 and increasing $L$ provides a better balance between representation space and compression efficiency. We refer to DDN with $K$=2 as Binary DDN (BinDDN). To our knowledge, BinDDN are the first generative model capable of directly transforming data into semantically meaningful binary strings. These binary strings represent a leaf node on a balanced binary tree.
@@
3.3 Techniques
Leak Choice. In each layer of the DDN, the output is conditioned on the image selected from the previous layer. This condition serves as a signal in the image domain, requiring the current layer to expend computational resources to extract features and interpret the choices made by the Sampler. However, for a DDN with a Single Shot Generator structure, the computation between two adjacent layers is minimal, involving only two convolutional layers. To facilitate faster understanding of the choices made by the Sampler in the subsequent layer, we have added extra convolutional layers to each output node. The features produced by these extra convolutions also serve as conditional inputs to the next layer. But these features don’t participate in the calculation of distance or loss in DDL.
@@ 3.3 Techniques
4 Experiments
We trained our models on a server equipped with 8 RTX2080Ti GPUs, setting the Chain Dropout probability to 0.05 by default. For the 64x64 resolution experiments, we utilized a DDN with 93M parameters, setting $K=512$ and $L=128$. In the CIFAR experiments, we employed a DDN with 74M parameters, setting $K=64$ and $L=64$. The MNIST experiments were conducted using a Recurrence Iteration Paradigm UNet model with 407K parameters, where $K=64$ and $L=10$. DDN is implemented on the foundation of the EDM Karras et al. (2022) codebase, with training parameters nearly identical to EDM. Both code and model weights shall be released for reproducing. More extended experiments exploring the properties of DDN can be found in the Appendix.
4.1 Qualitative and Quantitative Results
Generation Quality. fig. 6(a) and fig. 6(b) depict random generation results of DDN on CelebA-HQ-64x64 Karras et al. (2017) and FFHQ-64x64 Karras et al. (2019), verifying the model’s effectiveness in modeling facial data. The generation quality on CelebA-HQ appears superior to that on FFHQ, which is also reflected in the lower FID score (35.4 VS 43.1). We surmise this disparity arises from CelebA-HQ’s relatively cleaner backgrounds and less diverse facial data compared to FFHQ.
We conduct a qualitative comparison with a recent work, the Idempotent Generative Network (IGN) Shocher et al. (2024), accepted by ICLR 2024. Since IGN was only experimented on the CelebA dataset and did not release its code, our comparison is limited on the CelebA dataset. As depicted in fig. 7, our DDN demonstrate better generation capabilities over IGN.
(a) DDN (Ours), FID=35.4
(b) Idempotent Generative Network (IGN), FID=39
Figure 6(c) showcases the random generation results of DDN on the CIFAR dataset. In parallel, we present the FID score of DDN on unconditional CIFAR in table 1, comparing it against classical generative models. It is worth noting that modeling CIFAR remains a challenging task, especially for new and under-explored generative models like DDN. For instance, the recent work IGN Shocher et al. (2024) did not conduct experiments on CIFAR. Instead, we wish to emphasize the unique capability of DDN in “more general zero-shot conditional generatio”.
Conditional Training. Training a conditional DDN is quite straightforward, it only requires the input of the condition or features of the condition into the network, and the network will automatically learn $P(X|Y)$. fig. 6(d) and fig. 6(e) show the generation results of the class-conditional DDN on CIFAR and MNIST, respectively. The conditional DDN displays a good ability to learn the correct $P(X|Y)$ distribution.
Figure 8: Hierarchical Generation Visualization of DDN. We trained a DDN with output level $L=3$ and output nodes $K=8$ per level on MNIST dataset, its latent hierarchical structure is visualized as recursive grids. Each sample with a colored border represents an intermediate generation product. The samples within the surrounding grid of each colored-bordered sample are refined versions generated conditionally based on it (enclosed by the same color frontier). The small samples without colored borders are the final generated images. The larger the image, the earlier it is in the generation process, implying a coarse version. The large image in the middle is the average of all the generated images. More detailed visualization of $L=4$ will be presented in the Appendix.
Zero-Shot Conditional Generation. We trained a DDN model on the FFHQ-64x64 dataset and then tested the capability of the model for zero-shot conditional image generation using the CelebA-HQ-64x64 dataset, as shown in fig. 2. We presented the experiments of text-guided ZSCG using CLIP Radford et al. (2021) in the Appendix. Our generation process does not require gradient derivation or numerical optimization, nor does it need iterative steps. To the best of our knowledge, DDN is the first generative model that supports purely discriminative model as guide for zero-shot conditional generation.
In addition, we employed an off-the-shelf CIFAR classifier Phan (2021) to guide the generation of specific category images by a DDN model trained unconditionally on CIFAR. We want to emphasize that the classifier is an open-source, pre-trained ResNet18 model, with no additional modifications or retraining. fig. 9 displays images of various CIFAR categories generated by the model under the guidance of the classifier.
Latent analysis. We trained a DDN on the MNIST dataset with $K=8$ and $L=3$ to visualize both the hierarchical generative behavior of the DDN and the distribution of samples in the entire latent representation space, as shown in Figure 9.

### v3

@@ Abstract
We introduce a novel generative model, the Discrete Distribution Networks (DDN), that approximates data distribution using hierarchical discrete distributions. We posit that since the features within a network inherently capture distributional information, enabling the network to generate multiple samples simultaneously, rather than a single output, may offer an effective way to represent distributions. Therefore, DDN fits the target distribution, including continuous ones, by generating multiple discrete sample points. To capture finer details of the target data, DDN selects the output that is closest to the Ground Truth (GT) from the coarse results generated in the first layer. This selected output is then fed back into the network as a condition for the second layer, thereby generating new outputs more similar to the GT. As the number of DDN layers increases, the representational space of the outputs expands exponentially, and the generated samples become increasingly similar to the GT. This hierarchical output pattern of discrete distributions endows DDN with unique properties: more general zero-shot conditional generation and 1D latent representation. We demonstrate the efficacy of DDN and its intriguing properties through experiments on CIFAR-10 and FFHQ. The code is available at https://discrete-distribution-networks.github.io/
@@ 1 Introduction
To accurately fit the target distribution of large datasets, a substantial representational space is required. In the most extreme scenario, this space must be larger than the number of training data samples. However, current neural networks lack the feasibility to generate such a vast number of samples simultaneously. Therefore, we adopt a strategy from autoregressive models Van Den Oord et al. (2016) and partition this large space into a hierarchical conditional probability model. Each layer of this model needs only a small number of outputs. We then select one of these outputs as the output for that layer and use it as conditional input to the next layer. As a result, the output of the next layer will be more closely related to the selected conditional sample. If the number of layers is $L$ and the number of outputs per layer is $K$, then the output space of the network is $K^{L}$. Due to its exponential nature, this output space will be much larger than the number of samples in the dataset. Figure 1 shows how DDN generates images.
We posit that the contributions of this paper are as follows:
  • We introduce a novel generative model, termed Discrete Distribution Networks (DDN), which exhibit a more straightforward and streamlined principle and form.
  • For training the DDN, we propose the “Split-and-Prune” optimization algorithm, and a range of practical techniques.
  • We conduct preliminary experiments and analysis on the DDN, showcasing its intriguing properties and capabilities, such as zero-shot conditional generation and 1D latent representations.
@@ 2 Related Work
Deep Generative Model. Generative Adversarial Networks (GANs) Radford et al. (2016); Brock et al. (2019) and Variational Autoencoders (VAEs) Kingma & Welling (2014); van den Oord et al. (2018) are two early successful generative models. GANs reduce the divergence between the generated sample distribution and the target distribution through a game between the generator and discriminator. However, regular GANs cannot map samples back to the latent space, thus they cannot reconstruct samples. VAEs encode data into a simple distribution’s latent space through an Encoder, and the Decoder is trained to reconstruct the original data from this simple distribution’s latent space. Autoregressive models van den Oord et al. (2016) , with their simple principles and methods, model the target distribution by decomposing the target data into conditional probability distributions of each component. They can also compute the exact likelihood of target samples. However, the efficiency of these models is reduced when dealing with image data, which is not suitable to be decomposed into a sequence of components. Normalizing flow Kingma & Dhariwal (2018) is another class of generative models that can compute the likelihood. They use invertible networks to construct a mapping from samples to a noise space, and during the generation stage, they map back from noise to samples. Energy Based Models (EBM) Goyal et al. (2021); Song et al. (2021) with their high-quality and rich generative results, have led to the rise of diffusion models. However, their multi-step iterative generation process requires substantial computational resources. The recent Idempotent Generative Network Shocher et al. (2024) introduces a novel approach by training a neural network to be idempotent, mapping any input to the target distribution effectively.
Connections to VQ-VAE. While both VQ-VAE van den Oord et al. (2018) and DDN involve discrete representations, they differ significantly in their approach and capabilities. VQ-VAE enhances the traditional VAE by replacing the continuous latent space with discrete codebooks, thus achieving a more compact representation. VQ-VAE-2 Razavi et al. (2019) further improves this by employing a multi-scale hierarchical structure, thereby enhancing its representational power. However, the discrete representation in VQ-VAE remains two-dimensional, potentially leading to redundant information. Furthermore, VQ-VAE and its successors still rely on an additional prior network for generative modeling in the latent space. Notably, DDN can also serve as this prior model to effectively model the latent space of VQ-VAE. Other distinctions between DDN and VQ-VAE include the absence of an encoder and codebook in DDN, as well as its capacity for Zero-Shot Conditional Generation. VQ-VAEs are known to encounter codebook collapse, a problem that some researchers have addressed by reinitializing unused codes near frequently used ones Williams et al. (2020); Dhariwal et al. (2020). Our Split-and-Prune algorithm shares a similar core idea, albeit with some differences. While the reinitialization method aims to balance code usage to mitigate codebook collapse, our goal is to align the discrete distribution output by our network as closely as possible to the target distribution.
@@ 3 Discrete Distribution Networks
For computational efficiency, we adopted a decoder structure similar to the generator in GANs for coarse-to-fine image generation, as shown in fig. 3(b). We refer to this as the Single Shot Generator which is our default choice. As each layer of DDN can naturally input and output RGB domain data, DDN seamlessly support the Recurrence Iteration Paradigm fig. 3(c).
@@
How to generation. When the network performs the generation task, replacing the Guided Sampler with a random choice enables image generation. Given the exponential representational space of $K^{L}$ sample points and the limited number of samples in the training set, the probability of sampling an image with the same latent space as those in the training set is also exponentially low. For image reconstruction tasks, the process is almost identical to the training process, only substituting the training image with the target image to be reconstructed and omitting the L2 Loss part from the training process. The Final Generated Image in fig. 3(a) represents the final reconstruction result. The indices of the selected samples along the way form the target image’s latent $\mathbf{k}^{*}_{1:L}$, same as eq. 3. Therefore, the latent $\mathbf{k}^{*}_{1:L}$ is a sequence of integers with length $L$, which we regard as the hierarchical discrete representation of the target sample. The latent space exhibits a tree structure with $L$ layers and $K$ degrees per node, where each leaf node represents a sample point, and its latent denotes the indices of all nodes along the path to this leaf node, as shown in fig. 1(b). In the latent sequence, values placed earlier correspond to higher-level nodes in the tree, controlling the low-frequency information of the output sample, while later values tend to affect high-frequency information.
@@ 3.1 Optimization with Split-and-Prune
We have observed two primary issues resulting from each layer calculating loss only with the closest output samples to the ground truth (GT). Firstly, a problem similar to the“dead codebooks” in VQ-VAE, wherein output nodes that are not selected for a long time receive no gradient updates. During generation, these “dead nodes” are selected with equal probability, leading to poor output. The second issue is the probability density shift. For instance, in a one-dimensional asymmetric bimodal distribution, the target distribution is a mixture of two thin and tall Gaussian distribution functions, with one larger and one smaller peak. The means of these two peaks are -1 and 1, respectively. Therefore, half of the output samples with initial values less than 0 will be matched with samples from the larger peak and optimized towards the larger Gaussian distribution. Meanwhile, the other half of the output samples with values greater than 0 will be optimized towards the smaller Gaussian. A problem arises as the large and small peaks carry different probability masses but occupy equal portions of the output samples, resulting in the same sampling probability during generation.
@@
Inspired by the theory of evolution and genetic algorithms Katoch et al. (2021), we propose the Split-and-Prune algorithm to address the above issues, as outlined in algorithm 1. The Split operation targets nodes frequently matched by training samples, while the Prune operation addresses the issue of “dead nodes”, similar to those in Williams et al. (2020); Dhariwal et al. (2020). These nodes are akin to species in evolution, subject to diversification and extinction. During training, we track the frequency with which each node is matched by training samples. For nodes with excessive frequency, we execute the Split operation, cloning the node into two new nodes, each inheriting half of the old node’s frequency. Although these two new nodes have identical parameters and outputs, the next matched training sample will only be associated with one node. Therefore, the loss and gradient only affect one node’s parameters. Consequently, their parameters and outputs exhibit slight differences, dividing the old node’s match space into two. In subsequent training, the outputs of the two new nodes will move towards the centers of their respective spaces under the influence of the loss, diverging to produce more diverse outputs. For nodes with low matching frequency (dead nodes), we implement the Prune operation, removing them outright. Figure 4 illustrates the process of the Split-and-Prune operation and how it reduces the distance between the discrete distribution represented by the Output Nodes and the target distribution. The efficacy of the Split-and-Prune optimization algorithm is validated through examples of fitting 2D density maps in fig. 17.
@@ 3.2 Applications
Zero-Shot Conditional Generation (ZSCG). The DDN sampling process occurs directly within the sample space, which inherently facilitates Zero-Shot Conditional Generation. Each layer of the DDN generates multiple target samples, with a selected sample being forwarded as a condition to the next layer. This enables the generation of new samples in the desired direction, ultimately producing a sample that meets the given condition. In fact, the reconstruction process shown in fig. 1(a) is a ZSCG process guided by a target image. It is important to note that the target image is never directly input into the network. Instead, it shapes the generation outcomes by steering the sampling process.
To implement ZSCG, we replace the Guided Sampler in fig. 3(a) with a Conditional Guided Sampler. For instance, when generating an image of class $y_{i}$ guided by a classifier $g_{cls}$, we replace the “L2 Distance to Training Image” in the Guided Sampler of fig. 3(a) with the probability of each output image belonging to class $y_{i}$ according to the classifier. We then replace “argmin” with “argmax” to construct the Guided Sampler for this classifier. Similar to eq. 1, the sampling method is:
@@
After performing $L$ steps guided sampling and $L\times K$ steps of classification, the ZSCG result can be obtained without any gradient.
For super-resolution and colorization tasks, we construct a transform that converts the generated images into the target domain (low-resolution or grayscale). This approach significantly reduces the impact of the missing signal from the condition on the generated images, enabling DDN to successfully perform super-resolution tasks even when the source image has a resolution as low as $4\times 4$.
The use of “argmax” in the Guided Sampler causes each layer to select a fixed sample, resulting in the same image output under identical conditions, similar to greedy sampling in GPT. To increase diversity, we use Top-k sampling with $k=2$ for most ZSCG tasks, which balances diversity and condition appropriateness within the large generation space ($2^{L}$).
The versatility of ZSCG can be further enhanced by combining different Guided Samplers. For example, an image can guide the primary structure while text guides the attributes. The influence of each condition can be adjusted by setting their respective weights. Experiments involving the combination of different Guided Samplers are presented in appendix B.
@@
In our experiments, we set $K$=512 as the default, considering the balance between generation performance and training efficiency. However, from the perspective of data compression, setting $K$ to 2 and increasing $L$ provides a better balance between representation space and compression efficiency. We refer to DDN with $K$ = 2 as Taiji-DDN because of its similarity to the concept of Taiji in ancient Chinese philosophy, as described in fig. 18. To our knowledge, Taiji-DDN is the first generative model capable of directly transforming data into a semantically meaningful binary string which represents a leaf node on a balanced binary tree.
3.3 Techniques
Leak Choice. In each DDN layer, the output is conditioned on the image selected from the previous layer. This condition serves as a signal in the image domain, requiring the current layer to expend computational resources to extract features and interpret the choices made by the sampler. However, the computation between two adjacent layers in Single Shot Generator Paradigm is very small, involving only two convolutional layers. To facilitate a faster understanding of the choices made by the Sampler in the subsequent layer, we have added extra convolutional layers to each output node. The features produced by these extra convolutions also serve as conditional inputs to the next layer. But these features don’t participate in the calculation of distance or loss in DDL.
@@ 3.3 Techniques
4 Experiments
We trained our models on a server equipped with 8 RTX2080Ti GPUs, setting the Chain Dropout probability to 0.05 by default. For the 64x64 resolution experiments, we utilized a DDN with 93M parameters, setting $K=512$ and $L=128$. In the CIFAR experiments, we employed a DDN with 74M parameters, setting $K=64$ and $L=64$. The MNIST experiments were conducted using a Recurrence Iteration Paradigm UNet model with 407K parameters, where $K=64$ and $L=10$. DDN is implemented on the foundation of the EDM Karras et al. (2022) codebase, with training parameters nearly identical to EDM. More extended experiments exploring the properties of DDN can be found in the appendix.
4.1 Qualitative and Quantitative Results
Generation Quality. Figures 5(a) and 5(b) depict random generation results of DDN on CelebA-HQ-64x64 Karras et al. (2017) and FFHQ-64x64 Karras et al. (2019), verifying the model’s effectiveness in modeling facial data. The generation quality on CelebA-HQ appears superior to that on FFHQ, which is also reflected in the lower FID score (35.4 VS 43.1). We surmise this disparity arises from CelebA-HQ’s relatively cleaner backgrounds and less diverse facial data compared to FFHQ.
Table 1: Quantitative comparison on CIFAR-10. The data for VQ-VAE comes from Vuong et al. (2023). Data for other baselines comes from Bond-Taylor et al. (2021).
Method Type FID↓
DCGAN GAN 37.1
IGEBM EBM 38.0
VAE VAE 106.7
VQ-VAE VAE 117.4
Gated PixelCNN AR 65.9
GLOW Flow 46.0
DDN(ours) DDN 52.0
Figure 5(c) showcases the random generation results of DDN on the CIFAR dataset. We also present the FID score of DDN on unconditional CIFAR in table 1, comparing it with classical generative models. It is worth noting that modeling CIFAR remains a challenging task, especially for new and under-explored generative models like DDN. For instance, the recent work IGN Shocher et al. (2024) did not conduct experiments on CIFAR.
Conditional Training. Training a conditional DDN is quite straightforward, it only requires the input of the condition or features of the condition into the network, and the network will automatically learn $P(X|Y)$. Figures 5(d) and 5(e) show the generation results of the class-conditional DDN on CIFAR and MNIST, respectively. Figure 8 further demonstrates the combination of conditional generation and ZSCG on image-to-image tasks.
@@ 4.1 Qualitative and Quantitative Results
Zero-Shot Conditional Generation. We trained a DDN model on the FFHQ-64x64 dataset and then evaluated its zero-shot conditional image generation capability using the CelebA-HQ-64x64 dataset, as shown in fig. 2. In fig. 16, we presented the experiments of text-guided ZSCG using CLIP Radford et al. (2021). In particular, our generation process does not require gradient derivation, numerical optimization, or iterative steps. To the best of our knowledge, DDN is the first generative model to support the use of a purely discriminative model as a guide for zero-shot conditional generation.
In addition, we employed an off-the-shelf CIFAR classifier Phan (2021) to guide the generation of specific category images by a DDN model trained unconditionally on CIFAR. We want to emphasize that the classifier is an open-source, pre-trained ResNet18 model, with no additional modifications or retraining. Figure 7 displays images of various CIFAR categories generated by the model under the guidance of the classifier.
Latent analysis. We trained a DDN on the MNIST dataset with $K=8$ and $L=3$ to visualize both the hierarchical generative behavior of the DDN and the distribution of samples in the entire latent representation space, as shown in fig. 7.